OpenStudy (anonymous):

State the domain and range for the function, find the inverse, f^(-1)(x), and state the domain and range of the inverse. 1) y = 2x +1 2) y = 2 - 1/x

OpenStudy (cwrw238):

1) f)x) = 2x + 1 domain = all real numbers , range = same f^(-1)(x) = ( x - 1) / 2 domain and range is same as f(x)